OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y - x^2 = -3x - 10?

OpenStudy (fools101):

Try this web site an let me know if this helps you! http://www.mathway.com if not let me know ~Jess~

OpenStudy (anonymous):

Does it explain the process of solving quadratic equations? If it does ii will love you forever!

OpenStudy (anonymous):

It didnt work..... all i got was something weird.....

OpenStudy (fools101):

Yes it should if u sign up to it, I think.... okay if that don't work try this one: http://www.wolframalpha.com if this don't work let me know I try to fine other one for you!

OpenStudy (fools101):

hey is it suppose to be y-X^2 = 3X-10???

OpenStudy (anonymous):

oh ya sorry

hero (hero):

Process of Solving Quadratic Equations: 0. Given x^2 = -3x - 10 1. Write Equation in standard form: x^2 + 3x + 10 = 0 2. Find two numbers that add to get b, but multiply to get ac: In this case, 3 = 5 - 2; 10 = 5 x 2: 3. Rewrite b as the addition of those two numbers: x^2 + (5 - 2)x + 10 = 0 4. Distribute x across b: x^2 + 5x - 2x + 10 = 0 5. Factor the first two terms; then factor the last two terms: x(x + 5)-2(x + 5) = 0 6. Factor out the common term (in this case x + 5): (x+5)(x-2) = 0 7. Rewrite the equation using zero product property: x + 5 = 0 x - 2 = 0 8. Solve each equation for x: x = -5 x = 2 9. Finished

OpenStudy (anonymous):

Thank you so much hero!

hero (hero):

Warning, sometimes it might be difficult to find the two numbers